# Demystifying Workload Imbalances in Large Transformer Model Training over Variable-length Sequences

### v1

@@ 8.1 Experimental Setup
Figure 12: (a) Latency over 100 training iterations, with colors indicating the strategy selected for each iteration. (b) Comparison of load balancing between Hydraulis and Megatron-LM. We present the maximum and minimum propagation latency among all ranks for 5 iterations.
8.2 End-to-end Experiments
We first evaluate Hydraulis against three baselines across various configurations (baselines’ strategies are detailed in Appendix F). The results in Figure 10 demonstrate that Hydraulis achieves 1.32-2.66$\times$ improvement over the best baselines.
As DeepSpeed and Megatron-LM are limited to a single fixed strategy, they must adopt less efficient strategies to accommodate the memory consumption of long sequences, degrading the overall efficiency given most sequences are relatively short. Although HotSPa is capable of using multiple homogeneous strategies within an iteration to train sequences of varying lengths, it requires multiple times of model hot switching. This incurs significant overheads, especially when per-iteration time is short (e.g., in the “GitHub 7B 8GPUs F1” scenario, the average iteration time is merely 6.7s, while nearly 2s is spent on hot switching). On the contrary, Hydraulis handles variable-length sequences using dynamic heterogeneous strategies, while eliminating the overhead of strategy transition via disaggregating optimization and propagation.
@@ 8.2 End-to-end Experiments
8.3 Scalability
@@
We then assess the scalability of different systems by examining three factors: (a) number of GPUs, (b) maximum sequence length, and (c) tokens per iteration. Using the LLaMA 7B model and the CommonCrawl dataset, we start with a default setup of 16 GPUs, a maximum sequence length of 32K, and 200K tokens per iteration. By systematically altering one factor while keeping the others constant, we gain insights into each factor’s impact. The results are presented in Figure 11.
@@
Number of GPUs. As the number of GPUs increases, all systems experience a sub-linear reduction in latency due to the communication overhead introduced by parallelism. Specifically, for HotSPa, the communication overhead from hot switching does not decrease with the number of GPUs, causing its overall performance to be surpassed by Megatron-LM when using 64 GPUs. In contrast, Hydraulis introduces minimal additional communication overhead (only the LNK operators), which can overlap with computation. Consequently, its latency consistently outperforms other baselines.
@@
Maximum sequence length. As the maximum sequence length grows, DeepSpeed and Megatron-LM resort to less efficient strategies, resulting in a significant increase in latency. Notably, when Megatron-LM scales from 64K to 128K, it must enable cross-node TP, which greatly increases latency. In contrast, by customizing parallel strategies for sequences of different lengths, HotSPa and Hydraulis are less sensitive to maximum sequence length changes, with Hydraulis demonstrating superior performance over HotSPa.
@@
Tokens per iteration. As the number of tokens per iteration increases, all systems experience an almost linear increase in latency. However, HotSPa and Hydraulis exhibit a slightly lower growth rate, as they take account of the long-tail distribution of sequence lengths within iterations.
8.4 Performance Interpretation
@@
Figure 13: Ablation study. Error bars indicate the standard deviation across iterations. Total throughput changes as different components of Hydraulis are incrementally integrated.
@@
Case study. Figure 12(a) shows the optimal strategies selected by Hydraulis, along with the corresponding latencies, and summarizes the frequency of different strategies over 100 iterations. It can be seen that: (a) the actual propagation latency closely matches our estimation in most cases, and (b) variations across iterations lead to different strategy choices.
@@
Additionally, we examine the load balancing. Figure 12(b) shows that Megatron-LM, with fixed-length packing and even dispatching of packed sequences, results in a latency difference of 0.36-0.65$\times$ between the slowest and fastest ranks. In contrast, Hydraulis reduces this gap to just 0.02-0.10$\times$, verifying the significance of the two-stage sequence assignment.
Ablation study. Figure 13 presents an ablation study by progressively enabling different techniques in Hydraulis to evaluate their contributions: (a) Starting with a static, homogeneous parallel strategy, we dispatch sequences of varying lengths according to Eq. (3) and pack them to the maximum length. (b) Instead of packing sequences to the maximum length, we apply packing based on the results of Eq. (1). (c) We replace the original homogeneous parallel strategy with a heterogeneous strategy that supports the maximum sequence length. (d) Finally, we enable dynamic heterogeneous strategies, utilizing all Hydraulis designs.
The results show consistent improvements across (a)-(d), demonstrating that each component of the system plays a critical role in enhancing overall performance.
Latency breakdown. We further investigate the effectiveness of optimization-propagation disaggregation compared to conventional model states sharding, as well as the benefits brought by subgraphs scheduling. As shown in Table 1, we first adopt a fixed homogeneous parallel strategy (with the two-stage sequence assignment enabled), which utilizes the conventional all-gather and reduce-scatter interaction patterns (Figure 2(a)). Then, we employ dynamic heterogeneous strategies, introducing additional LNK operators (Figure 2(b)). Besides, we experiment both with and without subgraphs scheduling to measure the latency improvements from overlapping pull and push interactions with computation.
@@
P 6.78s 4.29s (1.58×1.58$\times$1.58 ×) 4.29s (1.58×1.58$\times$1.58 ×)
O 0.02s 0.02s (1.00×1.$0$) 0.02s (1.00×1.$0$)
O-P 1.11s 1.71s (0.69×0.69$\times$0.69 ×) 1.18s (0.94×0.94$\times$0.94 ×)
E2E 7.91s 6.02s (1.31×1.31$\times$1.31 ×) 5.49s (1.44×1.$4$)
Table 1: Breakdown of latency per iteration for different training strategies. "P" denotes propagation (Figure 2(b) step 3), "O" refers to optimization (Figure 2(b) step 6), and "O-P" represents the remaining overhead from the interaction between the two phases (Figure 2(b) steps 1-2 and 4-5). "E2E" stands for end-to-end training latency.
@@
It can be observed that the use of dynamic heterogeneous strategies improves the performance of propagation. However, the interaction between optimization and propagation becomes slower due to the additional communication overhead introduced by LNK operators. Despite this, the overhead remains relatively small and can be efficiently reduced by overlapping with computation, leading to only a minimal increase in latency. Moreover, this additional latency does not scale with the number of tokens per iteration, demonstrating the viability of disaggregating optimization and propagation.
9 Related Work
@@
Optimizations of distributed training. Numerous studies aim to enhance the efficiency of distributed training. In the realm of parallel methods, advancements in pipeline parallel focus on minimizing bubble time through enhanced scheduling [69, 49, 39, 85], while tensor parallel has progressed by overlapping communication with computation [82, 37, 11, 40]. In the context of model states sharding, efforts have been made to balance memory usage and latency through finer-grained sharding [13], efficient communication [81, 57], and multi-level storage solutions [73, 90]. Notably, these approaches are compatible with ours. There are also efforts focusing on automatically searching for the optimal parallel strategies [91, 38, 58, 84]. However, they only consider homogeneous strategies and overlook the data imbalance issues.
@@
Resources disaggregation. The idea of disaggregating storage from compute resources has been well-studied in cloud computing [74, 31, 53] and utilized in many AI applications [3, 42]. Similar concepts have also gained attention in large Transformer model inference, where disaggregating the prefilling and decoding phases allows for better throughput [67, 34, 92]. In our system, we adopt a similar design philosophy by separating the storage-heavy optimization phase from the compute-intensive propagation phase, enabling flexible execution and transitions of parallel strategies.
@@
Intermediate representations of parallel strategies. Graph-based intermediate representations are widely utilized in deep learning systems [1, 50, 41, 32]. By analyzing and optimizing these graphs, models can discover enhanced partitioning strategies and parallel schemes. Hydraulis extends this concept by simultaneously deducing and expressing graphs for different strategies. Using subgraphs, it distinctly highlights the unique communication patterns of each strategy. Besides, any graph-based optimization techniques can be applied to Hydraulis, allowing for further refinement for each strategy.
10 Conclusion and Future Work
@@
In this work, we addressed the often-overlooked issues of data sampling and packing imbalances in existing training systems, and introduced Hydraulis, a novel system that leverages parallel computing and data management co-designs to optimize the training performance of large Transformer models. Empirical results show that Hydraulis can achieve up to a 2.66×2.$6$ increase in throughput compared to existing systems.
@@
As a possible future direction, we intend to expand our strategy space by integrating more parallel methods, as our system designs are fully compatible with more forms of parallelism. We also plan to tackle data imbalance challenges in diverse domains (e.g., in the video domain, the significant variation in sequence lengths has already attracted considerable attention [14, 55, 75]). We believe our work can be adapted to a broader range of scenarios and applications.
@@ Appendix A Data imbalance
Figure 14: Curve of maximum sequence length across iterations. We sample 1.6M tokens from the CommonCrawl dataset per iteration and track the longest sequence.
A.1 Larger Batch Size
@@ A.2 End-to-end Influence of Attention
To better demonstrate the impact of the quadratic time complexity of the attention mechanism on end-to-end training, we train the 7B model. Sequences are packed in mini-batches using the first-fit-decreasing bin packing algorithm [66], with a maximum length of 32K. Each mini-batch contains 100K tokens, and we manually ensure that they will form 4 packed sequences (if not, we just skip that iteration for this demonstration experiment). Figure 15 shows the results. Despite carefully controlling tokens and the number of packed sequences, training time still fluctuates by about 20% in both settings. This variation primarily stems from differences in the computational cost of the attention mechanism, while other computational operations and communication overhead remain relatively stable.
Appendix B Cost Model

### v2

@@ 8. Evaluation
Figure 15. (a) Latency over 100 training iterations, with colors indicating the strategy selected for each iteration. (b) Comparison of load balancing between Hydraulis and Megatron-LM. We present the maximum and minimum propagation latency among all ranks for 5 iterations.
8.1. Experimental Setup
Environment
Our experiments are conducted with 8 nodes, each containing 8 Nvidia A800 GPUs, with a total number of 64 GPUs. GPUs within a node are connected via NVLink with a bandwidth of 400 GB/s, and nodes are interconnected using Infini-Band with a bandwidth of 200 GB/s.
Baselines
We compare Hydraulis with Megatron-LM (Shoeybi et al., 2019; Narayanan et al., 2021b; Korthikanti et al., 2022), DeepSpeed (Rasley et al., 2020; Jacobs et al., 2023), and HotSPa (Ge et al., 2024). Both Megatron-LM and DeepSpeed employ a single, static homogeneous strategy throughout the training process. HotSPa uses heuristics to configure multiple homogeneous strategies and enables switching between them within every iteration. However, due to the overhead of switching, HotSPa is limited to setting 3-4 strategies. We enumerate all competitors’ parallel strategies and select the optimal one to maximize baseline performance (see Appendix G (hyd, 2025) for details).
Models and datasets
Most existing large models are based on the decoder-only Transformer architecture. We select the widely used LLaMA2 series models and conduct experiments with 7B, 13B and 32B model sizes (Touvron et al., 2023), using CommonCrawl (TIIUAE, [n. d.]) and GitHub (CodeParrot, [n. d.]), which are widely used open-source datasets and have served as the major training corpora for many prestigious Transformer models.
Worloads and metrics
By default, we sample 100K tokens per iteration [^6] We consider 100K tokens per iteration for our 64-GPU cluster, scaled proportionally from the 16M used in LLaMA with an 8192-GPU cluster (Dubey et al., 2024). with a 32K context length, and use the Adam optimizer (Kingma and Ba, 2015). We conduct 20 warm-up iterations and then summarize the throughput (tokens per second) or per-iteration latency over 100 iterations.
8.2. End-to-end Experiments
@@
As DeepSpeed and Megatron-LM are limited to a single fixed strategy, they must adopt less efficient strategies to accommodate the memory consumption of long sequences, degrading the overall efficiency given most sequences are relatively short. Although HotSPa is capable of using multiple homogeneous strategies within an iteration to train sequences of varying lengths, it requires multiple times of model hot switching. This incurs apparent overheads, especially when per-iteration time is short (e.g., in the “GitHub 7B 8GPUs F1” scenario, the average iteration time is merely 6.7s, while nearly 2s is spent on hot switching). [^7] A more detailed analysis of HotSPa’s switching overhead can be found in Appendix H (hyd, 2025). On the contrary, Hydraulis handles variable-length sequences using dynamic heterogeneous strategies, while eliminating the overhead of strategy transition via disaggregating optimization and propagation.
Moreover, none of the baselines consider fine-grained sequence management. This problem is particularly severe for HotSPa— in some cases, there are only 1-2 long sequences per iteration, yet HotSPa is forced to adopt an individual strategy with all GPUs to process such long sequences, leaving a few GPUs with no sequences to process (e.g., in the “CommonCrawl 32B 64GPUs F2” scenario, HotSPa adopts $\left<TP,PP,CP\right>\times DP=\left<16,1,1\right>\times 4$ to process sequences ranging 16K-32K in length, often resulting in 3/4 or 1/2 GPUs being completely unused). Instead, Hydraulis, with the two-stage sequence assignment, meticulously organizes the training workloads of different sequences for better performance, as we will evaluate more in-depth later in §8.4.
8.3. Scalability
@@
We then assess the scalability of different systems by examining three factors: (a) number of GPUs, (b) maximum sequence length, and (c) tokens per iteration. Using the LLaMA 7B model and the CommonCrawl dataset, we start with a default setup of 16 GPUs, 32K maximum sequence length, and 200K tokens per iteration. By systematically altering one factor while maintaining others, we gain insights into each factor’s impact. Results are presented in Figure 14. Additionally, we analyze how (a) data imbalance and (b) problem-solving cost scale with increasing training workloads, particularly when both the cluster size (number of GPUs) and batch size (tokens per iteration) grow proportionally.
Assessment: number of GPUs
@@
As the number of GPUs increases, all systems experience a sub-linear reduction in latency due to the communication overhead introduced by parallelism. Specifically, for HotSPa, the communication overhead from hot switching does not decrease with the number of GPUs, causing its overall performance to be surpassed by Megatron-LM when using 64 GPUs. In contrast, Hydraulis introduces minimal additional communication overhead (only the LNK operators), which can overlap with computation. Consequently, its latency consistently outperforms other baselines.
Assessment: maximum sequence length
@@
As the maximum sequence length grows, DeepSpeed and Megatron-LM resort to less efficient strategies, resulting in a significant increase in latency. Notably, when Megatron-LM scales from 64K to 128K, it must enable cross-node TP, which greatly increases latency. In contrast, by customizing parallel strategies for sequences of different lengths, HotSPa and Hydraulis are less sensitive to maximum sequence length changes, with Hydraulis demonstrating superior performance over HotSPa.
Assessment: tokens per iteration
@@
As the number of tokens per iteration increases, all systems experience an almost linear increase in latency. However, HotSPa and Hydraulis exhibit a slightly lower growth rate, as they take account of the long-tail distribution of sequence lengths within iterations.
Analysis: data imbalance
When scaled to more GPUs, each iteration typically samples more training data, which may mitigate inter-iteration imbalance because longer sequences become more likely to be selected. To assess this, we conduct a 1024-GPU simulation with 16M tokens sampled per iteration (detailed in Appendix A.1 (hyd, 2025)). The results show that the maximum sequence length fluctuates between 16K-32K, indicating that inter-iteration imbalance still remains. Meanwhile, intra-iteration imbalance persists due to the dataset’s inherent sequence length distribution, while intra- and inter-pipeline imbalances originate from the discrepancy between the attention mechanism’s quadratic time complexity and linear memory complexity relative to sequence length. Since these imbalances are scale-independent, they persist at larger scales.
Analysis: problem-solving cost
We also conduct simulation experiments to evaluate the scalability of problem solving when scaled to 1024 GPUs (detailed in Appendix F (hyd, 2025)), which show that the problem-solving cost is well-bounded. To be specific, the strategy proposal (§7) incurs acceptable overhead (¡ 2 minutes) as it executes only once prior to training. Meanwhile, both sequence packing (§6.1) and dispatching (§6.2) operate on CPUs and can be fully overlapped with GPU training (i.e., the problem-solving for subsequent iterations occurs concurrently with current iteration training). Furthermore, larger cluster sizes provide additional CPU resources to increase the concurrency of problem solving, allowing us to further amortize the solving time per iteration.
@@
Figure 16. Ablation study. Error bars indicate the standard deviation across iterations. Total throughput changes as different components of Hydraulis are incrementally integrated.
8.4. Performance Interpretation
@@
Case study
Figure 15(a) shows the optimal strategies selected by Hydraulis, along with the corresponding latencies, and summarizes the frequency of different strategies over 100 iterations. It can be seen that: (a) the actual propagation latency closely matches our estimation in most cases, and (b) variations across iterations lead to different strategy choices.
To understand why using different strategy choices contributes, we compare the estimated latencies of two strategies, $S_{0}$ and $S_{3}$, over different iterations. While $S_{3}$ achieves $1.6\times$ speedup than $S_{0}$ in most cases due to its heterogeneous design, its maximum supportable sequence length is shorter than $S_{3}$, making it inapplicable for certain iterations. Besides, when the sampled batch contains fewer short sequences, it’s hard for $S_{3}$ to strike a balance (i.e., many sequences can only be processed by the pipeline with a TP degree of 8 while the other pipelines receive very few sequences), making it less efficient than $S_{0}$. These findings highlight the advantages of heterogeneous strategies, as well as the trade-off between latency and maximum sequence length across different strategies, demonstrating the benefits of supporting dynamic heterogeneous strategies for overall optimal performance.
@@
Ablation study
Figure 16 presents an ablation study by progressively enabling different techniques in Hydraulis to evaluate their contributions: (a) Starting with a static, homogeneous parallel strategy, we dispatch sequences of varying lengths according to Eq. (3) and pack them to the maximum length. (b) Instead of packing sequences to the maximum length, we apply packing based on the results of Eq. (1). (c) We replace the original homogeneous parallel strategy with a heterogeneous strategy that supports the maximum sequence length. (d) Finally, we enable dynamic heterogeneous strategies, utilizing all Hydraulis designs.
The results demonstrate that, on the one hand, when no optimizations are applied (i.e., Hydraulis-Base), our performance is on par with Megatron-LM, confirming the fairness of the comparison. On the other hand, consistent improvements are observed across (a)-(d), highlighting the critical role each system component plays in boosting overall performance.
Table 1. Breakdown of latency per iteration for different training strategies. “P” denotes propagation (Figure 3(b) step 3), “O” refers to optimization (Figure 3(b) step 6), and “O-P” represents the remaining overhead from the interaction between the two phases (Figure 3(b) steps 1-2 and 4-5). “E2E” stands for end-to-end training latency.
@@
P 6.78s 4.29s ($1.58\times$) 4.29s ($1.58\times$)
O 0.02s 0.02s ($1.00\times$) 0.02s ($1.00\times$)
O-P 1.11s 1.71s ($0.69\times$) 1.18s ($0.94\times$)
E2E 7.91s 6.02s ($1.31\times$) 5.49s ($1.44\times$)
Latency breakdown
We further investigate the effectiveness of optimization-propagation disaggregation compared to conventional model states sharding, as well as the benefits brought by subgraphs scheduling. As shown in Table 1, we first adopt a fixed homogeneous parallel strategy (with the two-stage sequence assignment enabled), which utilizes the conventional all-gather and reduce-scatter interaction patterns (Figure 3(a)). Then, we employ dynamic heterogeneous strategies, introducing additional LNK operators (Figure 3(b)). Besides, we experiment both with and without subgraphs scheduling to measure the latency improvements from overlapping pull and push interactions with computation.
@@
It can be observed that the use of dynamic heterogeneous strategies improves the performance of propagation. However, the interaction between optimization and propagation becomes slower due to the additional communication overhead introduced by LNK. Despite this, the overhead remains relatively small and can be efficiently reduced by overlapping with computation, leading to only a minimal increase in latency. Moreover, this additional latency does not scale with the number of tokens per iteration, demonstrating the viability of disaggregating optimization and propagation.
9. Related Work
Optimizations of distributed training
@@ 9. Related Work
Numerous studies aim to enhance the efficiency of distributed training. In the realm of parallel methods, advancements in pipeline parallel focus on minimizing bubble time through enhanced scheduling (Qi et al., 2023; Lamy-Poirier, 2023; Jiang et al., 2024a; Wu et al., 2024), while tensor parallel has progressed by overlapping communication with computation (Wang et al., 2023b; Jangda et al., 2022; Chen et al., 2024c; Jiang et al., 2024b). In the context of model states sharding, efforts have been made to balance memory usage and latency through finer-grained sharding (Chen et al., 2024a), efficient communication (Wang et al., 2023a; Maurya et al., 2024), and multi-level storage solutions (Ren et al., 2021; Zhao et al., 2025). Notably, these approaches are compatible with ours. There are also efforts focusing on automatically searching for the optimal parallel strategies (Zheng et al., 2022; Jia et al., 2019; Miao et al., 2022; Wang et al., 2024a; Li et al., 2025a; Wang et al., 2025b). However, they only consider homogeneous strategies and overlook the data imbalance issues.
Resources disaggregation
@@
The idea of disaggregating storage from compute resources has been well-studied in cloud computing (Shan et al., 2018; Guo et al., 2023; Lim et al., 2012) and utilized in many AI applications (Audibert et al., 2023; Jin et al., 2024). Similar concepts have also gained attention in large Transformer model inference, where disaggregating the prefilling and decoding phases allows for better throughput (Patel et al., 2024; Hu et al., 2024; Zhong et al., 2024). In our system, we adopt a similar design philosophy by separating the storage-heavy optimization phase from the compute-intensive propagation phase, enabling flexible execution and transitions of parallel strategies.
Intermediate representations of parallel strategies
@@
Graph-based intermediate representations are widely utilized in deep learning systems (Abadi et al., 2016; Lattner et al., 2021; Jin et al., 2020; Hagedorn et al., 2023). By analyzing and optimizing the graphs, models can discover enhanced and specialized parallel strategies. Hydraulis extends this by simultaneously deducing and expressing graphs for different strategies. Using subgraphs, it distinctly highlights the unique communication patterns of each strategy. Besides, any graph-based optimization techniques can be applied to Hydraulis, allowing for further refinement for each strategy.
10. Conclusion and Future Work
@@
In this work, we addressed the often-overlooked issues of data sampling and packing imbalances in existing training systems, and introduced Hydraulis, a novel system that leverages parallel computing and data management co-designs to optimize the training performance of large Transformer models. We proposed training with dynamic heterogeneous parallel strategies to address the data sampling imbalance and a two-stage sequence assignment to tackle the data packing imbalance. Empirical results show that Hydraulis can achieve up to a $2.66\times$ increase in throughput compared to state-of-the-art training systems.
@@
As a possible future direction, we intend to expand our strategy space by integrating more parallel methods (e.g., expert parallel (Lepikhin et al., 2020; Liu et al., 2025)), as our system designs are fully compatible with more forms of parallelism. We also plan to tackle data imbalance challenges in diverse domains (e.g., in the video domain, the significant variation in sequence lengths has already attracted considerable attention (Chen et al., 2024b; Liu et al., 2024; Shang et al., 2024; Xia et al., 2025)). We believe our work can be adapted to a broader range of scenarios and applications.
Acknowledgements.
This work is supported by National Natural Science Foundation of China (U23B2048, 62402011), PKU-Tencent joint research Lab, and High-performance Computing Platform of Peking University. Fangcheng Fu and Bin Cui are the corresponding authors.
@@ Appendix A Data Imbalance
Figure 17. Curve of maximum sequence length across iterations. We sample 1.6M tokens from the CommonCrawl dataset per iteration and track the longest sequence.
A.1. Larger Batch Size
@@ A.2. End-to-end Influence of Attention
To better demonstrate the impact of the quadratic time complexity of the attention mechanism on end-to-end training, we train the 7B model. Sequences are packed in mini-batches using the first-fit-decreasing bin packing algorithm (NVIDIA, 2023d), with a maximum length of 32K. Each mini-batch contains 100K tokens, and we manually ensure that they will form 4 packed sequences (if not, we just skip that iteration for this demonstration experiment). Figure 18 shows the results. Despite carefully controlling tokens and the number of packed sequences, training time still fluctuates by about 20% in both settings. This variation primarily stems from differences in the computational cost of the attention mechanism, while other computational operations and communication overhead remain relatively stable.
Appendix B Communication in O-P Disaggregation
B.1. Communication Bound
In this section, we provide formal proof of the communication bound in O-P disaggregation.
We consider $N$ GPUs and a model parameter of size $W$, and analyze the communication volume when each GPU participates as either a receiver or a sender during the pull operation, denoted as $V_{\text{recv}}$ and $V_{\text{send}}$, respectively.
For the receiver side analysis, consider a GPU belonging to a pipeline that adopts a parallel scheme $P=\langle TP,PP,CP\rangle$ during the propagation phase. The GPU must receive the complete parameter partition corresponding to its pipeline stage. Specifically:
  • If the parameter resides in the GPU’s current pipeline stage, it needs to receive a parameter partition of size $\frac{W}{TP}$.
  • If the parameter does not belong to the current stage, the receiving volume becomes 0 as no pull operation is required.
Therefore, we have:
$V_{\text{recv}}=\frac{W}{TP}\quad\text{or}\quad V_{\text{recv}}=0$
For the sender side analysis, our fully-sharded partitioning strategy during the optimization phase ensures each GPU stores exactly one parameter slice of size $\frac{W}{N}$. When $M$ GPUs (where $M\leq N$) request this particular parameter slice during propagation, the sending GPU’s communication volume is given by:
$V_{\text{send}}=\frac{W}{N}\times M$
Based on this, we can directly derive the worst-case communication scenario, which occurs when the parallel strategy takes the form $S=\langle 1,1,C\rangle\times\frac{N}{C}$, where $C$ is any positive integer divisor of $N$. This configuration represents the case where:
  • The system contains no tensor or pipeline parallelism ($TP=1$, $PP=1$).
  • Every GPU requires all parameter slices ($M=N$).
Consequently, each GPU’s communication volume reaches its theoretical maximum:
$V_{\text{recv}}=\frac{W}{1}=W\quad\text{and}\quad V_{\text{send}}=\frac{W}{N}\times N=W$
This worst-case scenario exactly corresponds to training with a homogeneous fully-sharded ZeRO-1 strategy, where the communication cost per GPU is bounded by $W$ and becomes independent of the total GPU count $N$.
An analogous analysis applies to the push operation.
B.2. Computation-Communication Overlap
In this section, we analyze the condition for overlapping communication with computation based on the derived communication bound.
For a matrix multiplication operation under the pipeline with parallel scheme $\langle\mathit{TP},\mathit{PP},\mathit{CP}\rangle$, consider input $X$ with shape $t/\mathit{CP},h]$ and weight matrix $W$ with shape $h,h^{\prime}/\mathit{TP}]$, where $t$ denotes the number of tokens, $h$ and $h^{\prime}$ represent hidden state dimensions. During the computation of this matrix multiplication, we aim to pull the parameter for the next matrix multiplication. As established in Appendix B.1, the maximum amount of data communicated would be $h\times h^{\prime}$ (i.e., the whole parameter size). Assuming half-precision (e.g., BF16) is used, the total communication volume is:
$\text{Comm}_{\text{bytes}}=2hh^{\prime}$
And the computational workload (FLOPs) for the matrix multiply (counting multiply-add as 2 operations) is:
$\text{FLOPs}=2\cdot\underbrace{\left(\frac{t}{\mathit{CP}}\right)}_{\text{rows}}\cdot\underbrace{h}_{\text{mid-dim}}\cdot\underbrace{\left(\frac{h^{\prime}}{\mathit{TP}}\right)}_{\text{cols}}=\frac{2thh^{\prime}}{\mathit{CP}\cdot\mathit{TP}}$
Let $F$ denote the GPU’s half-precision compute capability (FLOPS) and $B$ the interconnect bandwidth (bytes/sec). The computation time $T_{\text{comp}}$ and communication time $T_{\text{comm}}$ are:
$T_{\text{comp}}$ $=\frac{\text{FLOPs}}{F}=\frac{2thh^{\prime}}{\mathit{CP}\cdot\mathit{TP}\cdot F}$
$T_{\text{comm}}$ $=\frac{\text{Comm}_{\text{bytes}}}{B}=\frac{2hh^{\prime}}{B}$
To overlap communication with computation, we require:
$T_{\text{comp}}\geq T_{\text{comm}}$
Substituting the expressions and simplifying:
$\frac{2thh^{\prime}}{\mathit{CP}\cdot\mathit{TP}\cdot F}\geq\frac{2hh^{\prime}}{B}$
Canceling common terms ($2hh^{\prime}\neq 0$):
$\frac{t}{\mathit{CP}\cdot\mathit{TP}\cdot F}\geq\frac{1}{B}$
Solving for $t$:
$t\geq\mathit{TP}\cdot\mathit{CP}\cdot\frac{F}{B}$
This indicates that overlapping requires the token count $t$ to exceed a threshold determined solely by the parallel scheme ($\mathit{TP}$, $\mathit{CP}$) and hardware ratio ($F/B$), independent of parameter dimensions $h$, $h^{\prime}$.
As a concrete example, consider an A800 cluster running the 32B model with a pipeline configured with $\left<16,1,1\right$>, corresponding to Figure 15(a). The hardware specifications are (assuming the worst-case scenario where all communication goes through InfiniBand):
$F$ $=19.5\times 10^{12}\text{ FLOPS}$
$B$ $=200\times 10^{9}\text{ bytes/sec (worst-case)}$
The token threshold is then:
$t_{\text{threshold}}$ $=\mathit{TP}\cdot\mathit{CP}\cdot\frac{F}{B}=16\times 1\times\frac{19.5\times 10^{12}}{200\times 10^{9}}=1560$
This means that, theoretically, sequence lengths exceeding 1.5K after packing are sufficient to enable overlap (note that both GPU compute power and communication bandwidth may vary in real-world settings, so adjustments will be made based on empirical measurements). In practice, this condition is fully achievable in our setup.
Appendix C Cost Model